OpenStudy (anonymous):

What number must you add to complete the square? x^2+12x=-5

jimthompson5910 (jim_thompson5910):

take half of the x coefficient to get 12/2 = 6 then square this to get 6^2 = 36

jimthompson5910 (jim_thompson5910):

so you must add 36 to both sides

OpenStudy (anonymous):

72???

OpenStudy (anonymous):

the answer is 36 huh

jimthompson5910 (jim_thompson5910):

how are you getting 72?

OpenStudy (anonymous):

i thought u had to add 36+36=72

OpenStudy (anonymous):

x^2 + 12x + 5 = 0 x^2 + 12x = -5 Now, halve the linear term (12), square it and add it to both sides x^2 + 12 + 36 = -5 + 36 factor on the left, gather terms on the right (X + 6)2 = 31 (X + 6)2 - 31 = 0 ==============vertex form.... (- 6, - 31) ========================--------------vertex And that number added is 36

OpenStudy (anonymous):

so the answer is 36 right buddies

OpenStudy (anonymous):

yes